# E-values, Multiple Testing and Beyond

### v1

@@ 2.4 Connection between the BH and e-BH procedures
Let $\mathcal{S}_{\mathrm{BH}}$ be the set of rejections obtained through the BH procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (4). Then we have $\mathcal{S}_{\mathrm{BH}}=\mathcal{S}_{\mathrm{eBH}}$.
2.5 Connection between the BC and e-BH procedures
As noted in a recent work by Ren and Barber (2022), the BC procedure is equivalent to the e-BH procedure based on the following e-values:
$e_{i}=\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}},$
where $T_{\mathrm{BC}}$ is the threshold defined in (2).
3 Generalized BH and BC procedures
In this section, we extend the BH and BC procedures to allow a general class of rejection rules. This generalization enables the testing procedure to utilize cross-sectional information among the p-values and external structural information for each hypothesis, which often results in a higher multiple testing power. We shall establish the connection between these generalized procedures and the corresponding e-BH procedures.
3.1 Generalized BH procedure
We generalize the BH procedure to allow the rejection rule to take the form of $\varphi_{i}(p_{i})\leq t$, where $\varphi_{i}$ is a strictly increasing function and can differ for each $i$. Let us define $F_{i}(t)=\varphi_{i}^{-1}(t)$ as the inverse function of $\varphi_{i}$, and consider the rejection threshold
$T_{\text{GBH}}=\sup\left\{0<t\leq 1:\frac{\sum^{n}_{i=1}F_{i}(t)}{1\vee R(t)}\leq\alpha\right\},$ (5)
where $R(t)=\sum^{n}_{i=1}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$. The generalized BH (GBH) procedure rejects $H_{i}$ whenever $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}$. Similar to the BH procedure, the GBH procedure can be equivalently implemented in the following way. Define $\bar{F}(t)=n^{-1}\sum^{n}_{i=1}F_{i}(t)$ which is strictly increasing in $t$. We sort $q_{i}=\varphi_{i}(p_{i})$ in an ascending order, i.e., $q_{(1)}\leq\dots\leq q_{(n)}$ and find the largest $k$, represented as $\hat{k}$, for which $q_{(k)}\leq\bar{F}^{-1}\left(\frac{\alpha k}{n}\right).$ We then reject $H_{(i)}$ for all $i\leq\hat{k}$. The following proposition states that, under appropriate conditions in the form of $F_{i}$, the GBH procedure ensures FDR control.
Proposition 2.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. If $F_{i}(t)=c_{i}h(t)$ where $c_{i}$ is some constant and $h$ is a strictly increasing function of $t$, the GBH procedure controls the FDR at level $\alpha$.
As illustrated in the following examples, the GBH procedure aligns with several methods in the literature of structure adaptive multiple testing.
Example 1.
Let $\varphi_{i}(p)=p/\omega_{i}$, where $\omega_{i}$ denotes the weight for the $i$th hypothesis with $\omega_{i}>0$ and $\sum_{i=1}^{n}\omega_{i}=n$. The GBH procedure associated with this choice of $\varphi_{i}$ corresponds to the weighted BH procedure first introduced by Genovese et al. (2006). In this case, $F_{i}(t)=w_{i}t$ and the rejection threshold can be expressed as
$T_{\text{GBH}}=\sup\left\{0<t\leq 1\colon\frac{\sum_{i=1}^{n}w_{i}t}{1\vee R(t)}\leq\alpha\right\}=\sup\left\{0<t\leq 1\colon\frac{nt}{1\vee R(t)}\leq\alpha\right\},$
where $R(t)=\sum_{i=1}^{n}\mathbbm{1}\{p_{i}/w_{i}\leq t\}$.
Example 2.
Another choice of $\varphi_{i}$ is the local FDR (Efron, 2005; Sun and Cai, 2007) under a two-group mixture model. Specifically, we assume that the p-values are independently generated from the mixture density given by $\pi_{i}+(1-\pi_{i})(1-\kappa)p^{-\kappa}$, where $\pi_{i}\in(0,1)$ is the mixing proportion and $\kappa\in(0,1)$ controls the shape of the p-value density under the alternative. The rejection rule based on the local FDR under this model is given by
$\varphi_{i}(p)=\frac{\pi_{i}}{\pi_{i}+(1-\pi_{i})(1-\kappa)p^{-\kappa}}\leq t,$ (6)
which is the probability that the $i$th hypothesis is under the null given the observed p-value being $p$. It follows that
$F_{i}(t)=\varphi_{i}^{-1}(t)=\left(\frac{\pi_{i}(1-t)}{(1-\pi_{i})t(1-\kappa)}\right)^{-1/\kappa}=c_{i}h(t),$
where $c_{i}=\left\{\pi_{i}/\bigl{(}(1-\pi_{i})(1-\kappa)\bigr{)}\right\}^{-1/\kappa}$ and $h(t)=\left\{(1-t)/t\right\}^{-1/\kappa}$.
3.2 Connection between the GBH and e-BH procedures
Analogous to the BH procedure, we show that the GBH procedure is equivalent to the e-BH procedure applied to the following e-values:
$e_{i}=\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\text{GBH}}\}}{\sum^{n}_{i=1}F_{i}(T_{\mathrm{GBH}})},$ (7)
where $T_{\text{GBH}}$ is defined in (5). By the leave-one-out argument, we prove the following result.
Proposition 3.
Under the assumptions in Proposition 2, the e-values $e_{1},\dots,e_{n}$ defined in (7) satisfy Condition (3). As a consequence, the corresponding e-BH procedure controls the FDR at the desired level.
Additionally, we can prove that the e-BH procedure and the GBH procedure deliver the same set of rejections.
Theorem 2.
Let $\mathcal{S}_{\mathrm{GBH}}$ be the set of rejections obtained through the GBH procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (7). Then we have $\mathcal{S}_{\mathrm{GBH}}=\mathcal{S}_{\mathrm{eBH}}$.
3.3 Generalized BC procedure
In this section, we generalize the BC procedure with the rejection rule given by $\varphi_{i}(p_{i})\leq t$. We assume that the null p-value satisfies the following condition:
$P(p_{i}\leq a)\leq P(p_{i}\geq 1-a)=P(1-p_{i}\leq a),\quad\text{ for all }0\leq a\leq 0.5.$ (8)
Condition (8) is weaker than the mirror conservativeness in Lei and Fithian (2018), and it can be shown that super-uniformity implies (8). Indeed,
$P(1-p_{i}\leq a)\geq 1-P(p_{i}\leq 1-a)\geq 1-(1-a)=a\geq P(p_{i}\leq a).$
Assume that $\varphi_{i}$ is an increasing and continuous function, and define $F_{i}(x)=\sup\{0\leq p\leq 1\colon\varphi_{i}(p)\leq x\}$. We claim that
$P(\varphi_{i}(p_{i})\leq b)=P(p_{i}\leq F_{i}(b)).$
To see this, consider two cases. If $\varphi_{i}(p_{i})\leq b$, by the definition of $F_{i}(b)$, we have $p_{i}\leq F_{i}(b)$. On the other hand, if $p_{i}\leq F_{i}(b)$, then $\varphi_{i}(p_{i})\leq\varphi_{i}(F_{i}(b))=\lim_{p\uparrow F_{i}(b)}\varphi_{i}(p)\leq b$, where we use the fact that $\varphi_{i}$ is increasing to get the two inequalities, and the equality is due to the continuity of $\varphi_{i}$. Therefore, the above claim together with equation (8) implies that
$P(\varphi_{i}(p_{i})\leq b)=P(p_{i}\leq F_{i}(b))\leq P(1-p_{i}\leq F_{i}(b))=P(\varphi_{i}(1-p_{i})\leq b),$ (9)
for all $\varphi_{i}(0)\leq b\leq\varphi_{i}(0.5)$. Hence, we have
$\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$ $\approx\frac{\sum_{i\in\mathcal{H}_{0}}P(\varphi_{i}(p_{i})\leq t\})}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
$\leq\frac{\sum_{i\in\mathcal{H}_{0}}P(\varphi_{i}(1-p_{i})\leq t\})}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
$\approx\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
$\leq\frac{1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}},$
where the last term can be viewed as a conservative estimate of the FDP. Motivated by this observation, we define the threshold for the GBC procedure as
$T_{\text{GBC}}=\sup\left\{0<t\leq T_{\mathrm{up}}\colon\frac{1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}\leq\alpha\right\},$ (10)
which is the largest cutoff such that the FDP estimate is bounded above by $\alpha$, where $T_{\mathrm{up}}$ satisfies $T_{\mathrm{up}}<\min_{i}\varphi_{i}(0.5)$. The GBC procedure rejects $H_{i}$ whenever $\varphi_{i}(p_{i})\leq T_{\text{GBC}}$.
Proposition 4.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Assuming $\varphi_{i}$ is a monotonic increasing and continuous function for all $i$, then the GBC procedure ensures the FDR control at level $\alpha$.
Compared to the GBH procedure, the GBC approach affords us greater flexibility in selecting $\varphi_{i}$, as it no longer requires $\varphi_{i}$ to be strictly increasing and its (generalized) inverse function does not have to fulfill the condition in Proposition 2.
Example 3.
Suppose the p-value $p_{i}$ is generated independently from the two-group mixture model: $\pi_{i}f_{0}+(1-\pi_{i})f_{1,i}$, where $\pi_{i}\in(0,1)$ is the mixing proportion and $f_{0}$ and $f_{1,i}$ denote the p-value distributions under the null and alternative respectively. The local FDR is defined as
which is the posterior probability that the $i$th hypothesis is under the null given the observed p-value being $p$. The monotone likelihood ratio assumption (Sun and Cai, 2007) states that $f_{1,i}(p)/f_{0}(p)$ is decreasing in $p$. Under this assumption, $\varphi_{i}(p)=\mathrm{Lfdr}_{i}(p)$ is monotonically increasing in $p$ and thus fulfills the requirement in Proposition 4. Additionally, it has been shown in the literature that the rejection rule $\varphi_{i}(p_{i})=\mathrm{Lfdr}_{i}(p_{i})\leq t$ is optimal in the sense of maximizing the expected number of true positives among the decision rules that control the marginal FDR at level $\alpha$, see e.g., Sun and Cai (2007); Lei and Fithian (2018); Cao et al. (2022).
3.4 Connection between the GBC and e-BH procedures
We show that the GBC procedure is equivalent to the e-BH procedure with the following choice of e-values:
$e_{i}=\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\text{GBC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{\text{GBC}}\}},$ (11)
where $T_{\text{GBC}}$ is defined in (10). By equation (B.1) in the proof of Proposition 4, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n,$
which implies that the corresponding e-BH procedure controls the FDR at the desired level. Furthermore, the following theorem shows that the e-BH procedure with the e-values defined above is equivalent to the GBC procedure.
Theorem 3.
Let $\mathcal{S}_{\mathrm{GBC}}$ be the set of rejections obtained through the GBC procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (11). Then we have $\mathcal{S}_{\mathrm{GBC}}=\mathcal{S}_{\mathrm{eBH}}$.
4 Applications
Building upon the insights from the previous section, we shall develop new multiple testing procedures by aggregating the e-values resulting from different procedures (e.g., the BH and BC procedures) or the same procedure but applied to different subsets of the data. By showing that the aggregated e-values fulfill Condition (3), we have the e-BH procedure based on the aggregated e-values controls the FDR in finite sample. We illustrate this idea through three concrete applications below. The first application develops a hybrid procedure to leverage the strengths from the BH and BC procedures, aiming to improve their robustness and efficiency. The second application investigates fairness in the multiple testing context. In particular, given a sensitive/protected variable that partitions the data into several groups, we develop a fair multiple testing procedure that ensures the FDR within each group and the overall FDR across all groups are controlled at the same desired level simultaneously. Our last application proposes a new structure adaptive multiple testing procedure through cross-fitting and e-value aggregation. We demonstrate that the proposed procedure has guaranteed FDR control and can deliver comparable or even higher power than existing competitors.
4.1 A hybrid procedure
Empirical results in the literature suggest that neither the BH procedure nor the BC procedure dominates the other. Specifically, in the case of sparse signals (i.e., $n_{0}/n\approx 1$), the BH procedure typically has higher power than the BC procedure (Arias-Castro and Chen, 2017). In contrast, when the signal is dense, the BH procedure becomes conservative, making the BC procedure more preferable. However, in real-world applications, it is often impossible to know the signal strength and density in advance, and thus it is unclear which method would perform better. Applying both methods and reporting the results associated with the one that delivers more rejections does not ensure FDR control. Given the insights from Sections 2.4-2.5, we ask the question of whether it is possible to develop a new multiple testing procedure that has finite sample FDR control and is powerful across a wider range of signals by borrowing the strengths from both the BH and BC procedures. To address this question, we introduce a hybrid approach that utilizes e-values as a bridge to combine the BH and BC procedures, aiming for better robustness and efficiency. Let $e_{\mathrm{BH},i}$ and $e_{\mathrm{BC},i}$ be the e-values from the BH and BC procedures (at levels $\alpha_{\mathrm{BH}}$ and $\alpha_{\mathrm{BC}}$) respectively for testing the $i$th hypothesis. We define the weighted e-values
$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i},$
that aggregates the information from both the BH and BC procedures, and apply the e-BH procedure to the resulting e-values, where $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are non-negative weights. The detailed implementation of the hybrid procedure is described in Algorithm 1 below.
@@ 4.1 A hybrid procedure
1:p-values $p_{1},\dots,p_{n}$ and significance levels $\alpha_{\mathrm{BH}}$, $\alpha_{\mathrm{BC}}$ and $\alpha_{\mathrm{eBH}}$
2:Implement the BH procedure at the significance level $\alpha_{\mathrm{BH}}$. Compute the threshold $T_{\mathrm{BH}}$ using (1). Determine the e-value for the BH procedure with the formula:
@@
3:Implement the BC procedure at the significance level $\alpha_{\mathrm{BC}}$. Compute the threshold $T_{\mathrm{BC}}$ using (2). Determine the e-value for the BC procedure with the formula:
@@
4:Calculate the weighted averaged e-value by
$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i}.$ (12)
@@
5:Run the e-BH procedure utilizing the weighted averaged e-values at the significance level $\alpha_{\mathrm{eBH}}$.
6:The indices of rejected hypotheses.
4.1.1 Choice of significance levels and weights

### v2

@@ 3.1.1 Choosing Weights and FDR Control
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (6), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Then, the e-values specified in (4) with the weights defined via (5) satisfy Condition (2). Hence, the corresponding e-BH procedure controls the overall FDR in finite sample.
3.1.2 Illustrative Example of Data-Dependent Weights
In this subsection, we present a toy example to illustrate the effectiveness of data-dependent weights. Consider a scenario where we have two groups of p-values. The first group contains $n_{1}=100$ p-values, and the second group contains $n_{2}=1000$ p-values. We apply the BC method with a threshold of $0.05$. In each group, we identify $n_{a1}=n_{a2}=20$ significant features, yielding e-values of $e_{1}=100$ and $e_{2}=1000$ for the first and second groups, respectively.
Next, we combine the e-values from both groups into a single vector and sort them in decreasing order as $e_{(1)}\geq e_{(2)}\geq\cdots\geq e_{(n)}$, where $n=n_{1}+n_{2}=1100$. To select $40$ significant features, the combined e-values should satisfy $e_{(40)}\geq\frac{1100}{40\times 0.05}=550$. Similarly, to select $20$ significant features, the combined e-values should satisfy $e_{(20)}\geq\frac{1100}{20\times 0.05}=1100$. Without weighting, however, we find that $e_{(20)}=1000<1100$ and $e_{(40)}=100<550$. Thus, when the e-values are assembled without weights and the e-BH method is applied, no hypotheses are identified as significant.
Figure 1: Comparison of weighted and unweighted e-values. Circles represent weighted e-values, while triangles denote unweighted e-values. Different colors indicate two groups.
Figure 1 illustrates the scenario with only the non-zero e-values. Circles represent the weighted e-values, while triangles denote the e-values without weights. Different colors indicate the two groups. When e-values are combined without weights and the e-BH method is applied, no hypotheses are significant. In contrast, when the e-values are combined with weights, all hypotheses become significant. This demonstrates that data-dependent weights serve to increase the smaller e-values and decrease the larger e-values, thereby enhancing the method’s ability to identify more significant discoveries while maintaining control over the FDR.
3.1.3 Numerical Studies
We shall compare the finite sample performance of the proposed method with two naive approaches through simulations. The first method disregards the group information and directly applies the BC procedure to all p-values. We refer to this method as BC_Com for future reference. BC_Com has two shortcomings. Firstly, it may fail to control the group-wise FDRs, as illustrated in Setting E2. Secondly, it fails to ensure comparable power across different groups, resulting in the possibility of one group having high power. In contrast, the other group has nearly zero power, as illustrated in Setting E1. The same issue is also encountered by eBH_1. An alternative approach involves implementing the BC procedure for each group separately and combining all rejections. We call this method BC_Sep. Although BC_Sep effectively controls the FDR for individual groups, it does not guarantee the overall FDR control.
We first consider the case of two groups. To evaluate each method, we employ the following metrics: POW represents the overall power combining the rejections from both groups; POW1 denotes the power for the first group, while POW2 represents the power for the second group. Similarly, we can define FDR, FDR1, and FDR2. The empirical power and FDR are computed based on 1,000 independent Monte Carlo simulations.
In all settings, we assume that the p-values follow the uniform distribution on $[0,1]$ under the null. For the first group, the p-value is supposed to follow Beta($\alpha_{1},\beta_{1}$) under the alternatives, while for the second group, it follows Beta($\alpha_{2},\beta_{2}$) under the alternatives. The parameter values for different settings are detailed in Table D.1 in Supplement D.1.
Setting E1 corresponds to a scenario in which, for instance, the first group consists of ethnic minorities while the second group comprises ethnic majorities. The number of non-nulls is the same across the two groups. The alternative p-values in the first group are larger than those in the second group on average. The results for Setting E1 are presented in Table 1. BC_Com exhibits high power for the second group, yet its power in the first group is quite low. This is because the non-null p-values from the first group are not sufficiently small, and a combined analysis of the two groups demands a lower threshold, which thus fails to reject them. Additionally, BC_Sep has an inflated overall FDR in this case.
Setting E1 Setting E2
Method POW POW1 POW2 FDR FDR1 FDR2 POW POW1 POW2 FDR FDR1 FDR2
BC_Com 0.21 0.01 0.41 0.036 0.043 0.034 0.49 0.49 0.49 0.036 0.006 0.062
BC_Sep 0.378 0.336 0.420 0.060 0.035 0.035 0.416 0.727 0.105 0.056 0.048 0.017
eBH_1 0.075 0.000 0.149 0.021 0.000 0.021 0.024 0.000 0.049 0.010 0.000 0.010
eBH_2 0.127 0.128 0.126 0.012 0.013 0.010 0.079 0.082 0.077 0.009 0.005 0.012
eBH_Ada 0.212 0.185 0.238 0.027 0.019 0.019 0.289 0.499 0.079 0.038 0.034 0.013
Table 1: False discovery rate (FDR) and power for Settings E1 and E2 (nominal FDR level = 5%).
The results for Setting E2 are also presented in Table 1. We observe that BC_Com fails to control the FDR for the second group, which can be explained as follows. Due to the fact that the non-null p-values have a similar scale for both groups and the sample size of the first group is significantly smaller than that of the second group, BC_Com has a higher threshold compared to the BC procedure applied only to the second group. This can result in an FDR inflation in the second group for BC_Com. We also observe that BC_Sep suffers from an overall FDR inflation. In contrast, all three variants of the e-BH procedure control the group-wise and overall FDRs at the desired level. eBH_Ada has a much higher power than the other two e-value-based methods.
We present the results for $G=4$ in Supplement D.1, where we consider three different scenarios (Settings F1-F3). In particular, we note that BC_Com suffers from severe FDR inflation, with the empirical FDR reaching 0.318 at the 5% target level in Setting F2. In Setting F3, BC_Sep has an empirical overall FDR of 0.343, which is much higher than the 20% target level.
To summarize, as seen in Settings E2 and F2, BC_Com has no guarantee in controlling the group-wise FDR. On the other hand, BC_Sep fails to control the overall FDR, as observed in all the settings, particularly Settings E2 and F3. The e-BH-based approaches provide both group-wise and overall FDR control. However, eBH_1 and eBH_2 may suffer from power loss under certain scenarios. In contrast, eBH_Ada demonstrates consistent effectiveness across all settings by achieving both group-wise and overall FDR control, while maintaining reasonable power.
3.1.4 Real-Data Example
We illustrate the proposed method by conducting a differential abundance analysis using the microbiome dataset cdi_schubert, obtained from the MicrobiomeHD repository (Duvallet et al., 2017), originally collected in a case-control study comparing individuals with Clostridium difficile infection (CDI) to those without (nonCDI). After preprocessing, the feature table contains 2,293 operational taxonomic units (OTUs), representing bacterial taxa annotated at the phylum level. We then applied the LinDA method (Zhou et al., 2022) to identify taxa that differ between the CDI and nonCDI groups. Additional details about the dataset, the LinDA method, and implementation procedures are provided in Supplement D.2.
In this study, controlling the overall FDR ensures the reliability of global inference, whereas controlling the FDR within each phylum is essential for accurately interpreting results within biologically meaningful groups. The number of rejected taxa for each phylum is summarized in Table D.6 in Supplement D.2. The results show that for the phyla Bacteroidetes and Firmicutes, the eBH_Ada method yields fewer rejections compared to the BC_Com method, suggesting that BC_Com might inadequately control FDR within these specific groups. Conversely, for the phylum Proteobacteria, the eBH_Ada methods identify a greater number of rejections, mirroring the pattern observed in our simulation scenario E1, where the BC_Com method exhibits reduced power in certain groups. Moreover, the two data-independent weighting schemes, eBH_1 and eBH_2, produced no discoveries; consequently, we omit their results from the table. This outcome highlights the necessity of using data-dependent weights when aggregating e-values.
4 Aggregating E-Values From Different Results
In this section, we consider a scenario where we have $L$ sets of e-values, $\{e_{i}^{l}:i\in[n]\}_{l=1}^{L}$, potentially derived from $L$ distinct multiple testing procedures or the same multiple testing procedure with different test statistics or tuning parameters. Here, $\{e_{i}^{l}\}_{l=1}^{L}$ denotes the $L$ e-values associated with hypothesis $H_{i}$, and it satisfies $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{l}]\leq n.$ Our objective is to aggregate these $L$ sets of e-values into a single e-value vector $e_{1},\dots,e_{n}]$ that satisfies Condition (2).
We illustrate the proposed idea with a knockoff example in Supplement E, where we introduce a hybrid knockoff procedure that integrates multiple test statistics to achieve robustness across diverse modeling scenarios. Due to space constraints, a comprehensive discussion of the knockoff framework and the hybrid method is deferred to the Supplement.
@@ 4 Aggregating E-Values From Different Results
1:p-values $p_{1},\dots,p_{n}$ and significance levels $\alpha_{\mathrm{BH}}$, $\alpha_{\mathrm{BC}}$, and $\alpha_{\mathrm{eBH}}$
2:Execute the BH procedure at significance level $\alpha_{\mathrm{BH}}$. Compute the threshold $T_{\mathrm{BH}}$ using (1) with the BH-specified $m(t)$ and $R_{i}(t)$ defined in Table A.1. Calculate the e-value for the BH procedure as
@@
3:Execute the BC procedure at significance level $\alpha_{\mathrm{BC}}$. Compute the threshold $T_{\mathrm{BC}}$ using (1) with the BC-specified $m(t)$ and $R_{i}(t)$ defined in Table A.1. Calculate the e-value for the BC procedure as
@@
4:Compute the weighted averaged e-value using (7).
@@
5:Apply the e-BH procedure using the weighted average e-values at significance level $\alpha_{\mathrm{eBH}}$.
6:Indices of rejected hypotheses.
4.1 Hybrid Multiple-Testing Procedure
The second application of our proposed method leverages e-values to integrate results from both the BC and BH procedures. Empirical results in the literature indicate that neither the BH procedure nor the BC procedure consistently outperforms the other Arias-Castro and Chen (2017). In real-world applications, it is often impossible to determine which method will perform better. Applying both methods and reporting the results of the one that yields more rejections does not guarantee FDR control. In this section, we introduce a new multiple testing procedure that ensures finite-sample FDR control and maintains high power across a broader range of signals by leveraging the strengths of both the BH and BC procedures.
Let $e_{\mathrm{BH},i}$ and $e_{\mathrm{BC},i}$ denote the e-values from the BH and BC procedures (at significance levels $\alpha_{\mathrm{BH}}$ and $\alpha_{\mathrm{BC}}$, respectively) for testing the $i$th hypothesis. We define the weighted e-values as
$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i},$ (7)
which aggregates information from both the BH and BC procedures, where $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are non-negative weights. We then apply the e-BH procedure to these aggregated e-values to obtain our rejection set. The detailed implementation is given in Algorithm 2.
4.1.1 Choosing Significance Levels and Weights
We now discuss the choices of the significance levels $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ and the weights $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ in Algorithm 2, which play important roles in the hybrid procedure.